# Mixed-State Measurement-Induced Phase Transitions in Imaginary-Time Dynamics

### v2

@@ Abstract
Mixed-state phase transitions have recently attracted growing attention as a new frontier in nonequilibrium quantum matter and quantum information. In this work, we introduce the measurement-dressed imaginary-time evolution (MDITE) as a novel framework to explore mixed-state quantum phases and decoherence-driven criticality. In this setup, alternating imaginary-time evolution and projective measurements generate a competition between coherence-restoring dynamics and decoherence-inducing events. While reminiscent of monitored unitary circuits, MDITE fundamentally differs in that the physics is encoded in decoherent mixed states rather than in quantum trajectories. We demonstrate that this interplay gives rise to a novel class of mixed-state phase transitions, using numerical simulations of the one-dimensional transverse-field Ising model and the two-dimensional columnar dimerized Heisenberg model. Notably, the observed transitions do not fall into any previously established universality classes. Furthermore, we provide a diagrammatic representation of the evolving state, which naturally enables efficient studies of MDITE with quantum Monte Carlo and other many-body numerical methods, thereby extending investigations of mixed-state phase transitions to large-scale and higher-dimensional systems. In addition, the representation provides a natural interpretation of the phase transitions in terms of cluster formation within the simulations. Our results highlight MDITE as a powerful paradigm for investigating non-unitary dynamics and the fundamental role of decoherence in many-body quantum systems.
@@ I Introduction
Decoherence plays a pivotal role in quantum physics, explaining the emergence of classical behavior from quantum systems through their inevitable interaction with the environment [Zeh1970decoherence, Zurek198pointer, Zurek1982superselection, Zurek2001EnvInduced, Zurek2003review, Schlosshauer2005review, cai2013algebraic, cai2014identifying, yan2018interacting, Schlosshauer2019review]. By entangling with environmental degrees of freedom, quantum superpositions decohere into statistical mixtures within a stable pointer basis, suppressing quantum interference and giving rise to effectively classical outcomes [Zurek198pointer, Zurek1982superselection]. Beyond its fundamental implications, decoherence also poses a major challenge in quantum information science, as it undermines superposition and entanglement—essential resources for achieving quantum advantage [NielsenChuang2010, Bennett1996purification, Bravyi2005universal, Gottesman2009errorcorrection, Terhal2015errorcorrection].
@@
In contrast to uncontrollable environmental noise, projective quantum measurements provide a sharply defined and experimentally controllable way to introduce decoherence. This feature has made them a powerful tool for investigating open quantum systems. It has been recently discovered that, incorporating coherent dynamics into this setting leads to novel nonequilibrium phases and critical behaviors, collectively referred to as measurement-induced phase transitions (MIPTs) [liyaodong2018clifford, liyaodong2019clifford, SangSQ2021measure-protect, Skinner2019, Zabalo2022operator-scaling, fisher2023randomcircuit, Turkeshi2020measurement-criticality, Turkeshi2021from-infinite-to-zero-clicks, Lu2021spacetime-duality, Nahum2021measurement-entanglement, Sharma2022measurement-criticality, Szyniszewski2020universality-mpt, Tirrito2023fcs-mite, Zhou2023entanglement-criticality, Fux2024magiccircuit, Liushuo2024noise-mpt, Tsitsishvili2024mite-nonmarkovian, Chen2025subsystem, Chen2025mipt, zhu2023mipt]. The cannonical setting of MIPTs focuses on quantum trajectories [Wiseman1996quantum-trajectories, Plenio1998quantumjump, Daley2014quantumytrajectories]—ensembles of pure states conditioned on specific measurement outcomes—where the competition between scrambling unitary dynamics and disentangling measurements controls distinct scalings of entanglement.
@@
A complementary, yet fundamentally distinct, perspective arises by considering the measurement-averaged mixed state, which directly captures the universal characterstics of decoherence [Garratt2023nonlocal-measure, Nahum2020majoranadecoherence, Lee2023decoh_criticality, Amin2024symmetry-mixed-ent, Passarelli2024mixed-mite, Su2024higherform-symmetry, Tantivasadakarn2024longrange-ent-measure, Fanruihua2024diagnostics, wangzhu2025decoherence, guoyc2025lindladian, guoyc2025swssb, guoyc2025ldpo]. In contrast to conventional MIPTs which are fundamentally rooted in quantum trajectories of monitored unitary circuits, we instead consider a mixed-state perspective based on the interplay between imaginary-time evolution (ITE) and projective measurements—two intrinsically non-unitary processes governing the dynamics. Building on this idea,, we introduce the measurement-dressed imaginary-time evolution (MDITE) (Sec. II).
@@
Arising from a Wick rotation of the Schrödinger equation, ITE replaces real-time unitary operator $e^{-itH}$ with its non-unitary counterpart $e^{-\tau H}$, where $H$ is a Hamiltonian and $\tau$ denotes time. Unlike the oscillatory behavior of real-time quantum dynamics, ITE is intrinsically dissipative: high-energy excitations are exponentially suppressed so that generic initial states relax toward the ground state of the system described by $H$. Moreover, applying finite-time ITE to the maximally mixed state produces a thermal Gibbs state, highlighting its natural connection to finite-temperature statistical physics. While ITE was originally introduced as a mathematical tool for ground-state and thermal-state preparation, recent advances in quantum algorithms [Motta2020qite, McArdle2019vqite, Yuan2019theoryofvariational, Benedetti2021qite, Kondappan2023qite, Mao2023qite, Ding2024qite] and quantum hardware [Nishi2021qite-nisq, Cao2022qite, ZhangSX2024qite] have enabled its experimental realization. This also establishes the physical meaning of the MDITE we propose: on noisy quantum devices, the competition between ITE and decoherence arises naturally, providing a rich and realistic platform for studying novel many-body phenomena in open quantum systems.
@@
Specifically, our MDITE setup begins with a maximally mixed state and alternates between ITE controlled by a many-body Hamiltonian and random projective measurements. In the absence of measurement, ITE acts as a deterministic mechanism, steering the system toward the ground state of the Hamiltonian. Over extended times, this evolution suppresses classical uncertainty while reinforcing quantum coherence. In contrast, nonzero projective measurements tend to amplify classical uncertainty and induce decoherence, counteracting the ordering effects of ITE. This competition suggests the existence of mixed-state phase transitions with respect to the stationary state of the process driven by the strength of decoherence, evolution time, and parameters in the Hamiltonian. As we will demonstrate in Sec. IV and V, through large-scale numerical simulations of the one-dimensional transverse-field Ising model and the two-dimensional columnar dimerized Heisenberg model, we reveal the existence of such mixed-state phase transitions, which can be characterized using conventional linear order parameters—such as spin magnetization—without relying on information-theoretic proxies. Notably, our detailed analysis of the critical behavior shows that, to the best of our knowledge, these transitions do not fall into any previously established universality classes.
@@
It is important to emphasize that while our study focuses on a specific setup, the underlying framework is general and readily extends to other decoherence channels interacting with ITE. This establishes a broader paradigm of mixed-state phases and criticalities emerging from the interplay between imaginary-time dynamics and decoherence. Moreover, the rich diversity of ground states and exotic collective behaviors supported by different Hamiltonians significantly expands the scope for discovering novel many-body phenomena in open quantum systems. We will revisit the exciting potential for future exploration of this generalized paradigm in Sec. VI.
@@
This paper is organized as follows. In Sec. II, we formally define the MDITE protocol and highlight its key features. In Sec. III, we provide a schematic representation of the MDITE process, which not only clarify the setup but also facilitate numerical studies including quantum Monte Carlo (QMC) methods. Within this representation, the observed mixed-state phase transitions admit a natural interpretation in the picture of cluster formation in the QMC updates. The one- and two-dimensional Hamiltonians studied in this work, along with the corresponding numerical results for mixed-state phases and criticalities, are introduced and analyzed in Secs. IV and V. Finally, in Sec. VI, we summarize the work and discuss several promising directions for future research.
@@ II Protocol
In this section, we elaborate the measurement-dressed imaginary-time evolution (MDITE) protocol. While our discussion focuses on qubit systems, it can be readily generalized to systems with arbitrary local dimensions. As both the ITE and the projective measurements can be implemented on quantum circuits [Motta2020qite, McArdle2019vqite, Yuan2019theoryofvariational, Benedetti2021qite, Kondappan2023qite, Mao2023qite, Ding2024qite, Nishi2021qite-nisq, Cao2022qite, ZhangSX2024qite], we will describe the MDITE process in the language of quantum circuits.
@@
    Evolution: For each circuit layer $k=1,2,\dots$, the state evolves through two consecutive operations. First, a probabilistic projective measurement channel $\mathcal{M}_{p}$ is applied, where the measurement rate $p$ controls the probability that the qubits are measured. This is followed by an ITE with respect to $H$, i.e.,
    $\rho_{k}=\frac{1}{\mathcal{N}_{k}}e^{-\frac{\tau}{2}H}\mathcal{M}_{p}(\rho_{k-1})e^{-\frac{\tau}{2}H},$ (1)
    where $\mathcal{N}_{k}=\text{Tr}\!\left[e^{-\frac{\tau}{2}H}\mathcal{M}_{p}(\rho_{k-1})e^{-\frac{\tau}{2}H}\right]$ is the normalization factor ensuring $\text{Tr}(\rho_{k})=1$.
@@
    Output state: After applying $n_{d}$ layers, the final mixed state $\rho_{n_{d}}$ is obtained as the output.
In this study, we consider $\mathcal{M}_{p}$ to consist of probabilistic projective measurements in the computational basis. Specifically, each qubit is independently projectively measured with probability $p$, and left unchanged with probability (1p)(1-p). We are interested in the regime where, after a sufficiently large number of circuit layers $n_{d}$, the system approaches a stationary state, i.e., $\rho_{n_{d}}\approx\rho_{n_{d}+1}$. However, we stress that the existence of a stationary state is not guaranteed for general measurement channels. Nevertheless, for the examples considered in this work, such a stationary state indeed exists, as we will show in Secs. IV and V.
The dynamical process of MDITE involves a competition between two effects: the ITE operator $e^{-\frac{\tau}{2}H}$, which drives the system toward its ground state, and the measurement channel $\mathcal{M}_{p}$, which introduces decoherence and tends to produce a mixed state. As the computational basis is generally not the eigenbasis of $H$, ITE induces quantum fluctuations and reduces classical uncertainty. This suggests potential mixed-state phase transitions with respect to the stationary state $\rho_{n_{d}}$ when tuning the parameters in the protocol. It is important to note that the time step $\tau$ of ITE and the measurement rate $p$ serve as primary parameters controlling the behaviors of the stationary state $\rho_{n_{d}}$. Moreover, Hamiltonian parameters (e.g., coupling strength) also play a crucial role, resulting in a high-dimensional parameter space. This enriches the landscape of behaviors of the output state, rendering the setup of MDITE particularly compelling for investigating different novel many-body physics in mixed states.
@@ III.1 ITE without measurements
We begin by examining the case where no measurements are applied throughout the ITE processes, namely when the measurement rate is set to $p=0$. The state after the first circuit layer is
@@
Similarly, the output state after the second circuit layer, $\rho_{2}\propto e^{-2\tau H}$, is depicted in Fig. 1(b). We note that the intermediate states $|\ket{r^{\prime}}\bra{r}$ in the representation of $\rho_{2}$ [Fig. 1(b)] correspond to $|\ket{s^{\prime}}\bra{s}$ in the representation of $\rho_{1}$ [Fig. 1(a)], respectively. This follows from $\rho_{2}\propto e^{-\tau/2}\rho_{1}e^{-\tau/2}$, where the two ITE operators act on the upper and lower sides of $\rho_{1}$ in Fig. 1(a). In a manner analogous to tensor-network contraction, these operators propagate $\rho_{1}$, represented by $|\ket{r^{\prime}}\bra{r}$, into $\rho_{2}$, represented by $|\ket{s^{\prime}}\bra{s}$.
Notably, Fig. 1(a) and (b) can be interpreted as two ensembles, due to the summation over all possible $|\ket{s^{\prime}}\bra{s}$ in Eq. (3). We will revisit this notion of an ensemble in Sec. III.4. The cases of $\rho_{k}$ with $k>2$ follow in complete analogy, with their diagrammatic representations obtained recursively in the same way.
@@ III.2 ITE with deterministic measurements
We next consider the case in which all qubits are measured deterministically in each circuit layer. Note that the output state $\rho_{1}\propto e^{-\tau H}$ after the first circuit layer is independent of the choice of measurement channel $\mathcal{M}_{p}$ and whether $\mathcal{M}_{p}$ is applied, since measurements on a maximally mixed state always leave it unchanged. Therefore, the diagrammatic representation of $\rho_{1}$ under deterministic measurements remains Fig. 1(a).
In the second circuit layer, we first obtain the measurement-averaged state after the measurement channel $\mathcal{M}_{p}$, which is
@@
After the ITE, the resulting state is $\rho_{2}\propto e^{-\tfrac{\tau}{2}H}\,\bar{\rho}_{1}\,e^{-\tfrac{\tau}{2}H}$, which is the output state after the second circuit layer. Analogous to the discussion in Sec. III.1, $\rho_{2}$ can be represented diagrammatically as in Fig. 2(b). Here, the state $|\ket{r}\bra{r}$ corresponds to the state $|\ket{s}\bra{s}$ in $\bar{\rho}_{1}$ shown in Fig. 2(a). It is straightforward to generalize the discussion to a general $\rho_{k}$ $(k>2)$, and the diagrammatic representations also follow recursively.
Figure 2: Diagrammatic representations of the MDITE under deterministic projective measurements. (a) Compared with Fig. 1, the measurement on $\rho_{1}$ identifies the bra and ket components of $\rho_{1}$ in Eq. (3), forming a closed loop for the measurement-averaged state $\bar{\rho}_{1}$. (b) The output state $\rho_{2}$ after the second circuit layer.

### v3

@@ Abstract
Mixed-state phase transitions have recently attracted growing attention as a new frontier in nonequilibrium quantum matter and quantum information. In this work, we introduce the measurement-dressed imaginary-time evolution (MDITE) as a novel framework to explore mixed-state quantum phases and decoherence-driven criticality. In this setup, alternating imaginary-time evolution and projective measurements generate a competition between coherence-restoring dynamics and decoherence-inducing events. While reminiscent of monitored unitary circuits, MDITE fundamentally differs in that the physics is encoded in decoherent mixed states rather than in quantum trajectories. Using numerical simulations of the one-dimensional transverse-field Ising model and the two-dimensional columnar dimerized Heisenberg model, we demonstrate the existence of this kind of mixed-state phase transitions. Notably, these transitions appear to exhibit critical behavior inconsistent with known universality classes. In addition, we provide a diagrammatic representation of the evolving state, which naturally enables efficient studies of MDITE with quantum Monte Carlo and other many-body numerical methods, thereby extending investigations of mixed-state phase transitions to large-scale and higher-dimensional systems. Our results establish MDITE as a versatile platform for investigating mixed-state criticality and uncover new classes of decoherence-driven nonequilibrium phase transitions.
@@ I Introduction
Decoherence plays a pivotal role in quantum physics, explaining the emergence of classical behavior from quantum systems through their inevitable interaction with the environment [123, 131, 132, 71, 133, 86, 7, 8, 115, 87]. By entangling with environmental degrees of freedom, quantum superpositions decohere into statistical mixtures within a stable pointer basis, suppressing quantum interference and giving rise to effectively classical outcomes [131, 132]. Beyond its fundamental implications, decoherence also poses a major challenge in quantum information science, as it undermines superposition and entanglement—essential resources for achieving quantum advantage [68, 3, 5, 28, 98].
@@
In contrast to uncontrollable environmental noise, projective quantum measurements provide a sharply defined and experimentally controllable way to introduce decoherence. This feature has made them a powerful tool for investigating open quantum systems. It has been recently discovered that, incorporating coherent dynamics into this setting leads to novel nonequilibrium phases and critical behaviors, collectively referred to as measurement-induced phase transitions (MIPTs) [47, 48, 84, 89, 122, 25, 103, 102, 51, 65, 88, 94, 100, 119, 26, 49, 101, 11, 10, 129]. The cannonical setting of MIPTs focuses on quantum trajectories [111, 73, 16]—ensembles of pure states conditioned on specific measurement outcomes—where the competition between scrambling unitary dynamics and disentangling measurements controls distinct scalings of entanglement.
@@
A complementary, yet fundamentally distinct, perspective arises by considering the measurement-averaged mixed state, which directly captures the universal characterstics of decoherence [27, 66, 43, 63, 70, 90, 95, 24, 106, 30, 31, 32, 130]. In contrast to conventional MIPTs which are fundamentally rooted in quantum trajectories of monitored unitary circuits, we instead consider a mixed-state perspective based on the interplay between imaginary-time evolution (ITE) and projective measurements—two intrinsically non-unitary processes governing the dynamics. Building on this idea,, we introduce the notion of measurement-dressed imaginary-time evolution (MDITE) (Sec. II).
@@
Arising from a Wick rotation of the Schrödinger equation, ITE replaces real-time unitary operator $e^{-itH}$ with its non-unitary counterpart $e^{-\tau H}$, where $H$ is a Hamiltonian and $\tau$ denotes time. Unlike the oscillatory behavior of real-time quantum dynamics, ITE is intrinsically dissipative: high-energy excitations are exponentially suppressed so that generic initial states relax toward the ground state of the system described by $H$. Moreover, applying finite-time ITE to the maximally mixed state produces a thermal Gibbs state, highlighting its natural connection to finite-temperature statistical physics. While ITE was originally introduced as a mathematical tool for ground-state and thermal-state preparation, recent advances in quantum algorithms [64, 60, 121, 2, 41, 58, 20] and quantum hardwares [69, 9, 125] have enabled its experimental realization. This also establishes the physical meaning of the MDITE we propose: on noisy quantum devices, the competition between ITE and decoherence arises naturally, providing a rich and realistic platform for studying novel many-body phenomena in open quantum systems.
@@
Our MDITE setup begins with a maximally mixed state and alternates between ITE controlled by a local Hamiltonian and random projective measurements. In the absence of measurement, ITE acts as a deterministic mechanism, steering the system toward the ground state of the Hamiltonian. Over extended times, this evolution suppresses classical uncertainty while reinforcing quantum coherence. In contrast, nonzero projective measurements tend to amplify classical uncertainty and induce decoherence, counteracting the ordering effects of ITE. This competition suggests the existence of mixed-state phase transitions with respect to the steady state of the process driven by the strength of decoherence, evolution time, and parameters in the Hamiltonian.
As we will demonstrate in Secs. IV, V and VI, through large-scale numerical simulations of the one-dimensional transverse-field Ising model and the two-dimensional columnar dimerized Heisenberg model, we reveal the existence of such mixed-state MIPTs. Remarkably, these transitions can be characterized by conventional linear order parameters, such as spin magnetization, without relying on information-theoretic diagnostics typically used in MIPTs based on quantum trajectories [47, 48, 84, 89, 122, 25, 103, 102, 51, 65, 88, 94, 100, 119, 26, 49, 101, 11, 10, 129]. Supported by theoretical discussions and numerical analysis of the critical behavior, our results indicate that, to the best of our knowledge, these transitions do not belong to any previously established universality class.
@@
It is important to emphasize that while our study focuses on a specific setup, the underlying framework is general and readily extends to other decoherence channels interacting with ITE. This establishes a broader paradigm of mixed-state phases and criticalities emerging from the interplay between imaginary-time dynamics and decoherence. Moreover, the rich diversity of ground states and exotic collective behaviors supported by different Hamiltonians significantly expands the scope for discovering novel many-body phenomena in open quantum systems. We will revisit the exciting potential for future exploration of this generalized paradigm in Sec. VII.
@@
This paper is organized as follows. In Sec. II, we introduce the MDITE protocol and discuss its key features. In Sec. III, we present a diagrammatic representation of the MDITE process, which clarifies the setup and enables efficient numerical approaches, including quantum Monte Carlo (QMC). Within this framework, the mixed-state MIPTs admit a natural interpretation in terms of cluster formation in the QMC updates. The one- and two-dimensional Hamiltonians studied in this work, together with the corresponding results for mixed-state phases and critical behavior, are presented in Secs. IV and VI. In particular, Sec. V provides important theoretical insights by examining some continuous limits in which the dynamics map onto effective equilibrium problems in a doubled Hilbert space, thereby elucidating the underlying mechanism of the transitions we discovered qualitatively. Finally, in Sec. VII, we summarize our findings and outline several directions for future research.
@@ II Protocol
In this section, we formally define the measurement-dressed imaginary-time evolution (MDITE) protocol. While our discussion focuses on qubit systems, it can be readily generalized to systems with arbitrary local dimensions.
@@
    Evolution: For each discrete time step $k=1,2,\dots$, the state evolves through two consecutive operations:
    • A probabilistic projective measurement channel $\mathcal{E}_{p}$ in the computational basis is applied, where the measurement rate $p$ controls the probability that the qubits are measured;
    • An ITE with respect to $H$.
    The resulting state is
    $\rho_{k}=\frac{1}{\mathcal{N}_{k}}e^{-\frac{\tau}{2}H}\mathcal{E}_{p}[\rho_{k-1}]e^{-\frac{\tau}{2}H},$ (1)
    where $\mathcal{N}_{k}=\text{Tr}\!\left(e^{-\frac{\tau}{2}H}\mathcal{E}_{p}[\rho_{k-1}]e^{-\frac{\tau}{2}H}\right)$ is the normalization factor ensuring $\text{Tr}(\rho_{k})=1$.
@@
    Output state: After $n_{d}$ steps, the final mixed state $\rho_{n_{d}}$ is obtained as the output.
Specifically, we consider that each qubit is independently measured with probability $p$, and left unchanged with probability (1p)(1-p). Mathematically,
$\mathcal{E}_{p}=\prod_{i=1}^{N}\mathcal{E}_{p,i},$ (2)
with
$\mathcal{E}_{p,i}[\rho]\equiv$ $(1-p)\rho$ (3)
[…]
Here $|\ket{0}\equiv\ket{\uparrow}$ and $|\ket{1}\equiv\ket{\downarrow}$ denote the eigenstates of the local Pauli operator $Z_{i}$ acting on the $i$th qubit.
We are interested in the regime where, after a sufficiently large number of time steps $n_{d}$, the system approaches a steady state, i.e., $\rho_{n_{d}}\approx\rho_{n_{d}+1}\approx\rho_{\infty}$. However, we stress that the existence of $\rho_{\infty}$ is not guaranteed for general measurement channels. Nevertheless, for the examples considered in this work, such a steady state indeed exists, as we will show in Secs. IV and VI.
The dynamical process of MDITE involves a competition between two effects: the ITE operator $e^{-\frac{\tau}{2}H}$, which drives the system toward its ground state, and the measurement channel $\mathcal{E}_{p}$, which introduces decoherence and tends to produce a mixed state. As the computational basis is generally not the eigenbasis of $H$, ITE induces quantum fluctuations and reduces classical uncertainty. This suggests potential mixed-state phase transitions with respect to the steady state $\rho_{\infty}$ when tuning the parameters in the protocol.
It is important to note that the time length $\tau$ of each ITE and the measurement rate $p$ serve as primary parameters controlling the behaviors of the steady state $\rho_{\infty}$, as we will discuss in Sec. V. Moreover, Hamiltonian parameters (e.g., coupling strength) also play a crucial role, resulting in a high-dimensional parameter space. This enriches the landscape of behaviors of the output state, rendering the setup of MDITE particularly compelling for investigating and discovering novel many-body physics in open quantum systems.
@@ III Diagrammatic representation of the monitored imaginary-time evolution
Next, we introduce a diagrammatic representation to describe the evolving state $\rho_{n_{d}}$ in the MDITE process. This representation is powerful as it both enhances conceptual understanding and facilitates numerical studies, such as the efficient QMC method we introduce for the MDITE problem in Sec. III.5. Moreover, we note that analytical treatments of the MDITE steady state for general $\tau$ and $p$ are challenging, and only certain continuous limits are tractable, as we will discuss in Sec. V.
@@
We remark that this section is primarily technical. Readers mainly interested in the results and analysis may skip directly to Sec. IV without loss of continuity.
@@ III.1 ITE without measurements
We begin by examining the case where no measurements are applied throughout the ITE processes, namely when the measurement rate is set to $p=0$. The state after the first time step is
@@
Similarly, the output state after the second step, $\rho_{2}\propto e^{-2\tau H}$, is depicted in Fig. 1(b). We note that the intermediate states $|\ket{r^{\prime}}\bra{r}$ in the representation of $\rho_{2}$ [Fig. 1(b)] correspond to $|\ket{s^{\prime}}\bra{s}$ in the representation of $\rho_{1}$ [Fig. 1(a)], respectively. This follows from $\rho_{2}\propto e^{-\tau/2}\rho_{1}e^{-\tau/2}$, where the two ITE operators act on the upper and lower sides of $\rho_{1}$ in Fig. 1(a). In a manner analogous to tensor-network contraction, these operators propagate $\rho_{1}$, represented by $|\ket{r^{\prime}}\bra{r}$, into $\rho_{2}$, represented by $|\ket{s^{\prime}}\bra{s}$.
Notably, Fig. 1(a) and (b) can be interpreted as two ensembles, due to the summation over all possible $|\ket{s^{\prime}}\bra{s}$ in Eq. (5). The cases of $\rho_{k}$ with $k>2$ follow in complete analogy, with their diagrammatic representations obtained recursively in the same way.
@@ III.2 ITE with deterministic measurements
We next consider the case in which all qubits are measured deterministically in each time step. Note that the output state $\rho_{1}\propto e^{-\tau H}$ after the first time step is independent of the choice of measurement channel $\mathcal{E}_{p}$ and whether $\mathcal{E}_{p}$ is applied, since measurements on a maximally mixed state always leave it unchanged. Therefore, the diagrammatic representation of $\rho_{1}$ under deterministic measurements remains Fig. 1(a).
In the second time step, we first obtain the measurement-averaged state after the measurement channel $\mathcal{E}_{p}$, which is
@@
After the ITE, the resulting state is
$\rho_{2}\propto e^{-\tfrac{\tau}{2}H}\,\bar{\rho}_{1}\,e^{-\tfrac{\tau}{2}H},$ (7)
which is the output state after the second time step. Analogous to the discussion in Sec. III.1, $\rho_{2}$ can be represented diagrammatically as in Fig. 2(b). Here, the state $|\ket{r}\bra{r}$ corresponds to the state $|\ket{s}\bra{s}$ in $\bar{\rho}_{1}$ shown in Fig. 2(a). It is straightforward to generalize the discussion to a general $\rho_{k}$ $(k>2)$, and the diagrammatic representations also follow recursively.
Figure 2: Diagrammatic representations of the MDITE under deterministic projective measurements. (a) Compared with Fig. 1, the measurement on $\rho_{1}$ identifies the bra and ket components of $\rho_{1}$ in Eq. (5), forming a closed loop for the measurement-averaged state $\bar{\rho}_{1}$. (b) The output state $\rho_{2}$ after the second time step.